# Spatial Analysis for AI-segmented Histopathology Images: Methods and Implementation

### v1

@@ Abstract
Quantitatively characterizing the spatial organization of cells and their interaction is essential for understanding cancer progression and immune response. Recent advances in machine intelligence have enabled large-scale segmentation and classification of cell nuclei from digitized histopathology slides, generating massive point pattern and marked point pattern datasets. However, accessible tools for quantitative analysis of such complex cellular spatial organization remain limited. In this paper, we first review $27$ traditional spatial summary statistics, areal indices, and topological features applicable to point pattern data. Then, we introduce SASHIMI (Spatial Analysis for Segmented Histopathology Images using Machine Intelligence), a browser-based tool for real-time spatial analysis of artificial intelligence (AI)-segmented histopathology images. SASHIMI computes a comprehensive suite of mathematically grounded descriptors, including spatial statistics, proximity-based measures, grid-level similarity indices, spatial autocorrelation measures, and topological descriptors, to quantify cellular abundance and cell-cell interaction. Applied to two cancer datasets, oral potentially malignant disorders (OPMD) and non-small-cell lung cancer (NSCLC), SASHIMI identified multiple spatial features significantly associated with patient survival outcomes. SASHIMI provides an accessible and reproducible platform for single-cell-level spatial profiling of tumor morphological architecture, offering a robust framework for quantitative exploration of tissue organization across cancer types.
@@ 1 Introduction
Tissue morphology provides fundamental insights into organ function and disease mechanisms. In tumor tissues, the complex spatial arrangement of malignant and non-malignant cells forms the tumor microenvironment (TME) (hanahan2011hallmarks), which directly influences cancer progression, immune response, and therapeutic outcomes (zhai2017spatial; wang2023deep). Histopathology imaging remains the gold standard for visualizing cellular and tissue-level morphological architecture, offering rich histological detail on nuclei, lymphocytes, glands, and stromal structures, together providing “a comprehensive view of disease and its effect on tissue” (gurcan2009histopathological). Consequently, microscopic examination of hematoxylin and eosin (H&E)-stained slides is an essential steop in the diagnosis and staging of numerous diseases. Yet this workflow relies heavily on expert pathologists to detect and interpret subtle spatial and morphological patterns within highly complex images. The process is time-consuming, subjective, and prone to substantial inter and intra-observer variability (van2010interobserver; cooper2015novel).
@@
The rapid advancement of artificial intelligence (AI) has reshaped histopathology imaging analysis, enabling quantitative, scalable, and reproducible assessment of complex tissue structures (SALTZ2018181; corredor2019spatialb; wang2019convpath; sobhani2022spatiala; wang2022spatialc). Deep learning–based segmentation methods now delineate millions of individual cells from H&E–stained slides and automatically predict phenotypic or morphological labels that previously required manual pathologist annotation (wang2019convpath). As a result, characterizing cellular spatial organization and cell–cell interactions has become a central objective in modern computational pathology, though it remains challenging due to the inherent complexity and heterogeneity of tissue architecture. These AI-driven advances enable a shift from subjective, qualitative inspection to rigorous quantitative tissue modeling, allowing researchers to extract biologically and clinically meaningful imaging biomarkers and features.
@@
An AI-segmented histopathology image can be naturally represented as a marked point pattern, where each cell is characterized by its spatial coordinates and associated phenotypic or morphological label. This representation bridges digital pathology with spatial point pattern analysis and topological data analysis, providing a rigorous statistical framework for studying cellular spatial organization with the TME. A growing body of work shows that spatial correlation among cell types carry important prognostic value. For example, hierarchical Bayesian models have been used to quantify heterogeneous tumor–immune interactions, which are strongly associated with patient survival outcomes (li2019bayesiana; li2019bayesianb). More recent studies by moon2023using, zhang2023deep, stolz2024relational, and zhang2024bayesian demonstrate that geometric and topological descriptors capturing tissue shape can effectively stratify patient survival outcomes and characterize cellular and dynamical patterns. These findings underscore the clinical importance of quantifying multiscale cellular and architectural organization in tumor tissues through spatial and shape statistics.
@@
Despite these methodological advances, there is still no systematic review of spatial and shape statistics for histopathology image analysis, and, more importantly, no online, integrative tool that enables researchers to efficiently quantify and visualize cellular spatial organization across large-scale AI-segmented histopathology images. Recent efforts, such as spatial pathology features (SPF) (vu2022spf) and spatial quantification of pathology features (Spatial-QPFs) (li2024spatialqpfs), have demonstrated the clinical value of spatial summary statistics. However, these approaches rely on a limited set of spatial summary statistics and are available only as offline R packages, restricting usability and scalability.
@@
To address these limitations, we first provide a comprehensive summary of $27$ spatial summary statistics, areal indices, and topological features applicable to point pattern data. We then introduce SASHIMI (Spatial Analysis of Segmented Histopathology Images using Machine Intelligence), a unified web-based platform that integrates these spatial and topological methods into a scalable, accessible, and reproducible analytical framework. SASHIMI focuses on integration, standardization, and real-time implementation of diverse analytical approaches for characterizing cellular spatial organization within tumor tissues. The platform encompasses three complementary analytical modules: (i) spatial summary statistics, capturing intercellular proximity and interaction patterns; (ii) topological and graph-based features, describing connectivity and structural complexity; and (iii) areal similarity indices, quantifying spatial heterogeneity across tissue compartments. Implemented as a real-time, browser-based interface, SASHIMI enables researchers to perform comprehensive exploratory spatial analyses without requiring extensive programming expertise. By applying SASHIMI to two case studies: (i) the oral potentially malignant disorder (OPMD) cohort from the Erlotinib Prevention of Oral Cancer (EPOC) trial at The University of Texas MD Anderson Cancer Center and (ii) the non–small-cell lung cancer (NSCLC) cohort from the National Lung Screening Trial (NLST), we demonstrate that spatial features derived through this integrated platform are significantly associated with patient survival outcomes. These findings highlight SASHIMI’s value as a practical and reproducible tool for quantitative digital pathology.
@@
The remainder of this paper is organized as follows. Section 2 describes (marked) point pattern data extracted from histopathology images. Section 3 introduces the imaging feature extraction framework, covering spatial summary statistics, areal data indices, and topological features. Section 4 details the implementation of the SASHIMI online tool, including system design, input data requirements, output visualization modules, and user workflow. Section 5 presents two case studies, followed by a threshold-based significance analysis of the identified spatial biomarkers. Finally, Section 6 concludes the paper and discusses potential directions for future research.
@@ 3.1 Spatial Summary Statistics
Spatial summary statistics are functions that analyze the spatial characteristics of points based on their locations and relationships. These functions characterize properties such as clustering, regularity, and inter-type spatial relationships (baddeley2016spatial). Given marked point pattern data extracted from segmented pathology images, SASHIMI automatically computes the corresponding spatial summary functions and produces graphical visualizations. Since computed spatial features are functional data, SASHIMI displays the computed functions alongside the input image, providing an intuitive visual interpretation of spatial organization.
The tumor tissue exhibits spatial heterogeneity across multiple scales, where cells may cluster locally while showing regular spacing at larger distances, and different cell types may exhibit attraction, repulsion, or independence. To capture this complexity, SASHIMI computes a comprehensive suite of spatial summary statistics organized into function families, providing both univariate analyses of single cell types and bivariate analyses of inter-type interactions. For example, Ripley’s $K$-function (ripley1977modelling) quantifies the expected number of neighboring points within distance $r$ of a typical point, revealing clustering or dispersion patterns. The $G$-function characterizes the distribution of nearest-neighbor distances, while the $J$-function combines inter-point and empty-space information to detect deviations from complete spatial randomness.
All spatial summary statistics in SASHIMI are computed using the R package spatstat (baddeley2016spatial), which implements rigorous edge-correction methods to mitigate boundary effects in finite observation windows.
@@ 3.2.1 Spatial autocorrelation measures
Spatial autocorrelation measures quantify the degree of spatial dependence between neighboring quadrats. These include Moran’s I (moran1950notes), Geary’s C (geary1954contiguity), and the Quadrat Count Statistic, which assess whether quadrats with similar cell densities tend to cluster spatially or exhibit random or dispersed patterns.
@@ 3.2.2 Similarity indices
Similarity indices measure compositional overlap or dissimilarity between spatial distributions of different cell types. SASHIMI computes the Jaccard Index, the Morisita-Horn Index (morisita1962sigma), and cosine similarity to quantify the degree of co-localization among cell types within tissue regions.
@@
Computes the probability of two distributions from the same discrete domain $Q_{X}$, overlap. where:
@@
Together, these measures provide complementary perspectives on spatial organization: autocorrelation statistics capture global patterns of spatial dependence, while similarity indices describe local compositional relationships. All computed metrics are returned as scalar features in a single-row data frame. Computations are supported by the R packages spatstat (baddeley2016spatial) and spdep (bivand2015spdep).
@@ 3.3 Topological Features
Topology, a branch of mathematics that focuses on spaces that are invariant under continuous transformations, forms the basis of topological data analysis (TDA). It enables characterization of spatial patterns through topological features such as the number of connected components (dimension zero features) and loops (dimension one features). In SASHIMI, topological features of marked point pattern data are computed using persistent homology with the witness complex, a method that approximates the shape of the data by encoding its local connectivity structure (de2004topological). The resulting output is represented as a multiset of birth–death intervals. From these intervals, we extract summary statistics that quantitatively characterize the topological features across multiple spatial scales. For each dimension, we compute distributional statistics of birth times, death times, and feature lifetimes, including their minima, maxima, means, and standard deviations, along with the total number of detected features. This functional representation provides a complementary view of the spatial organization beyond classical distance-based descriptors. Computations are performed using the Python package GUDHI (gudhi:urm).

### v2

@@ Abstract
Quantitatively characterizing the spatial organization of cells and their interaction is essential for understanding cancer progression and immune response. Recent advances in machine intelligence have enabled large-scale segmentation and classification of cell nuclei from digitized histopathology slides, generating massive point pattern and marked point pattern datasets, which in turn create a growing need for accessible and standardized tools for downstream spatial statistical analysis. However, accessible, reproducible, and integrative tools for exploratory and feature-based quantitative analysis of such complex cellular spatial organization remain limited. In this paper, we do not aim to provide a comparative review or performance evaluation, but instead systematically organize and operationalize 27 widely used spatial summary statistics, areal indices, and topological features applicable to point pattern data within a unified computational framework. We introduce SASHIMI (Spatial Analysis for Segmented Histopathology Images using Machine Intelligence), a browser-based tool for real-time spatial analysis of artificial intelligence (AI)-segmented histopathology images. SASHIMI enables the computation of a comprehensive suite of mathematically grounded descriptors, including spatial statistics, proximity-based measures, grid-level similarity indices, spatial autocorrelation measures, and topological descriptors, to quantify cellular abundance and cell-cell interaction. Applied to two cancer datasets, oral potentially malignant disorders (OPMD) and non-small-cell lung cancer (NSCLC), SASHIMI is used to illustrate the types of spatial and architectural patterns that can be captured, with several extracted features showing associations with patient survival outcomes. SASHIMI provides an accessible and reproducible platform for single-cell-level spatial profiling of tumor morphological architecture, enabling interactive visualization and exploratory analysis of cellular spatial organization without requiring programming expertise.
@@ 1 Introduction
Tissue morphology provides fundamental insights into organ function and disease mechanisms. In tumor tissues, the complex spatial arrangement of malignant and non-malignant cells forms the tumor microenvironment (TME) (Hanahan and Weinberg, 2011), which directly influences cancer progression, immune response, and therapeutic outcomes (Zhai et al., 2017; Wang et al., 2023). Histopathology imaging remains the gold standard for visualizing cellular and tissue-level morphological architecture, offering rich histological detail on nuclei, lymphocytes, glands, and stromal structures, together providing “a comprehensive view of disease and its effect on tissue” (Gurcan et al., 2009). Consequently, microscopic examination of hematoxylin and eosin (H&E)-stained slides is an essential step in the diagnosis and staging of numerous diseases. Yet this workflow relies heavily on expert pathologists to detect and interpret subtle spatial and morphological patterns within highly complex images. The process is time-consuming, subjective, and prone to substantial inter and intra-observer variability (Van Den Bent, 2010; Cooper et al., 2015). As histopathology analysis increasingly shifts toward large-scale, cell-resolved digital workflows enabled by AI-based segmentation, there is a growing demand for quantitative and reproducible tools that enable a broader research community to characterize cellular spatial organization beyond expert-driven statistical programming environments.
@@
The rapid advancement of artificial intelligence (AI) has reshaped histopathology imaging analysis, enabling quantitative, scalable, and reproducible assessment of complex tissue structures (Saltz et al., 2018; Corredor et al., 2019; Wang et al., 2019; Sobhani et al., 2022; Wang et al., 2022). Deep learning–based segmentation methods now delineate millions of individual cells from H&E–stained slides and automatically predict phenotypic or morphological labels that previously required manual pathologist annotation (Wang et al., 2019). Importantly, these advances primarily operate at the image segmentation and cell classification stage. Once segmentation is completed, downstream analysis of cellular spatial organization and cell–cell interactions relies on statistical and mathematical modeling of the resulting point pattern data rather than deep learning models themselves. As a result, characterizing cellular spatial organization and cell–cell interactions has become a central objective in modern computational pathology, though it remains challenging due to the inherent complexity and heterogeneity of tissue architecture. This shift highlights a critical methodological gap between AI-driven image segmentation and statistically principled spatial analysis that remains accessible to end users.
@@
An AI-segmented histopathology image can be naturally represented as a marked point pattern, where each cell is characterized by its spatial coordinates and associated phenotypic or morphological label. This representation bridges digital pathology with spatial point pattern analysis and topological data analysis, providing a rigorous statistical framework for studying cellular spatial organization with the TME. A growing body of work shows that spatial correlation among cell types carries important prognostic value. For example, hierarchical Bayesian models have been used to quantify heterogeneous tumor–immune interactions, which are strongly associated with patient survival outcomes (Li et al., 2019d, b). Additional distance-based, co-localization, and functional summary approaches have also been proposed to capture spatial heterogeneity and interaction patterns in multiplex and histopathology imaging data, including quantile-based biomarkers, spatial co-localization indices, and comparative evaluation frameworks (Yi et al., 2023; Fridley et al., 2024; Soupir et al., 2025). More recent studies by Moon et al. (2023), Zhang et al. (2023), Stolz et al. (2024), and Zhang et al. (2024) demonstrate that geometric and topological descriptors capturing tissue shape can effectively stratify patient survival outcomes and characterize cellular and dynamical patterns. While these approaches focus on statistical modeling and inference, such as distance-matrix–based modeling, point process formulations, and functional ANOVA frameworks (Masotti et al., 2023; Seal et al., 2024; Mohammed et al., 2024), their practical application often requires extensive preprocessing, feature engineering, and custom implementations. In parallel, AI-assisted frameworks have been developed to extract morphological and shape-based features from segmented histopathology images. For example, SAFARI (Fernández et al., 2022) focuses on quantifying nuclear shape characteristics from AI-segmented images to support downstream statistical analysis. While such approaches advance morphology-aware image analysis, they primarily operate at the level of shape representation and feature extraction, leaving the systematic characterization of cellular spatial organization and cell–cell interactions as a distinct and complementary challenge. Taken together, these studies underscore the clinical importance of quantifying multiscale cellular and architectural organization in tumor tissues using complementary spatial and shape-based statistical descriptors.
@@
Despite these methodological advances, there remains a lack of integrative and accessible computational platforms that enable researchers to efficiently quantify, visualize and explore cellular spatial organization across large-scale AI-segmented histopathology images. Rather than the absence of individual methods, the primary limitation lies in the fragmentation of existing spatial and topological techniques across disparate software packages and analytical pipelines, as also highlighted in recent statistical studies on quantitative cancer imaging (Mohammed et al., 2024; Soupir et al., 2025). Recent efforts, such as spatial pathology features (SPF) (Vu et al., 2022) and spatial quantification of pathology features (Spatial-QPFs) (Li, 2024), have demonstrated the clinical value of spatial summary statistics. However, these approaches rely on a limited set of spatial summary statistics and are available primarily as offline R packages, which limits interactive exploration, standardized output, and rapid integration into downstream statistical workflows.
@@
To address these limitations, we focus on the design and implementation of a unified software platform that organizes, standardizes, and operationalizes widely used spatial analytical techniques for histopathology imaging data. Rather than providing a comparative evaluation of methods, we incorporate $27$ established spatial summary statistics, areal indices, and topological features applicable to point pattern data into a single computational framework. We introduce SASHIMI (Spatial Analysis of Segmented Histopathology Images using Machine Intelligence), a unified web-based platform that integrates these spatial and topological methods into a scalable, accessible, and reproducible analytical framework. SASHIMI focuses on integration, standardization, and real-time implementation of diverse analytical approaches for characterizing cellular spatial organization within tumor tissues. The platform encompasses three complementary analytical modules: (i) spatial summary statistics, capturing intercellular proximity and interaction patterns; (ii) topological and graph-based features, describing connectivity and structural complexity; and (iii) areal similarity indices, quantifying spatial heterogeneity across tissue compartments. Implemented as a real-time, browser-based interface, SASHIMI enables researchers to perform comprehensive exploratory spatial analyses without requiring extensive programming expertise. Through two illustrative case studies: oral potentially malignant disorders (OPMD) from the Erlotinib Prevention of Oral Cancer (EPOC) trial and non–small-cell lung cancer (NSCLC) from the National Lung Screening Trial (NLST), we demonstrate how SASHIMI facilitates exploratory investigation of cellular spatial architecture and the generation of clinically relevant hypotheses. Together, these examples underscore SASHIMI’s role as a practical and reproducible platform for quantitative digital pathology.
@@
The remainder of this paper is organized as follows. Section 2 describes (marked) point pattern data extracted from histopathology images. Section 3 introduces the imaging feature extraction framework, covering spatial summary statistics, areal data indices, and topological features. Section 4 details the implementation of the SASHIMI online tool, including system design, input data requirements, output visualization modules, and user workflow. Section 5 presents two illustrative case studies demonstrating the use of SASHIMI for exploratory spatial analysis, followed by a sensitivity analysis examining how exploratory findings vary across commonly used significance thresholds. Finally, Section 6 concludes the paper and discusses potential directions for future research.
@@ 3.1 Spatial Summary Statistics
Spatial summary statistics are functions that analyze the spatial characteristics of points based on their locations and relationships. These functions characterize properties such as clustering, regularity, and inter-type spatial relationships (Baddeley et al., 2016). Given marked point pattern data extracted from segmented pathology images, SASHIMI automatically computes the corresponding spatial summary functions and produces graphical visualizations. Since computed spatial features are functional data, SASHIMI displays the computed functions alongside the input image, providing an intuitive visual interpretation of spatial organization.
The tumor tissue exhibits spatial heterogeneity across multiple scales, where cells may cluster locally while showing regular spacing at larger distances, and different cell types may exhibit attraction, repulsion, or independence. To capture this complexity, SASHIMI computes a comprehensive suite of spatial summary statistics organized into function families, providing both univariate analyses of single cell types and bivariate analyses of inter-type interactions. For example, Ripley’s $K$-function (Ripley, 1977) quantifies the expected number of neighboring points within distance $r$ of a typical point, revealing clustering or dispersion patterns. The $G$-function characterizes the distribution of nearest-neighbor distances, while the $J$-function combines inter-point and empty-space information to detect deviations from complete spatial randomness.
All spatial summary statistics in SASHIMI are computed using the R package spatstat (Baddeley et al., 2016), which implements rigorous edge-correction methods to mitigate boundary effects in finite observation windows.
@@ 3.2.1 Spatial Autocorrelation Measures
Spatial autocorrelation measures quantify the degree of spatial dependence between neighboring quadrats. These include Moran’s I (Moran, 1950), Geary’s C (Geary, 1954), and the Quadrat Count Statistic, which assess whether quadrats with similar cell densities tend to cluster spatially or exhibit random or dispersed patterns.
@@ 3.2.2 Similarity Indices
Similarity indices measure compositional overlap or dissimilarity between spatial distributions of different cell types. SASHIMI computes the Jaccard Index, the Morisita-Horn Index (Morisita, 1962), and cosine similarity to quantify the degree of co-localization among cell types within tissue regions.
@@
where:
@@
Together, these measures provide complementary perspectives on spatial organization: autocorrelation statistics capture global patterns of spatial dependence, while similarity indices describe local compositional relationships. All computed metrics are returned as scalar features in a single-row data frame. Computations are supported by the R packages spatstat (Baddeley et al., 2016) and spdep (Bivand et al., 2015).
@@ 3.3 Topological Features
Topology, a branch of mathematics that focuses on spaces that are invariant under continuous transformations, forms the basis of topological data analysis (TDA). It enables characterization of spatial patterns through topological features such as the number of connected components (dimension zero features) and loops (dimension one features). In SASHIMI, topological features of marked point pattern data are computed using persistent homology with the witness complex, a method that approximates the shape of the data by encoding its local connectivity structure (De Silva and Carlsson, 2004). The resulting output is represented as a multiset of birth–death intervals. From these intervals, we extract summary statistics that quantitatively characterize the topological features across multiple spatial scales. For each dimension, we compute distributional statistics of birth times, death times, and feature lifetimes, including their minima, maxima, means, and standard deviations, along with the total number of detected features. This functional representation provides a complementary view of the spatial organization beyond classical distance-based descriptors. Computations are performed using the Python package GUDHI (The GUDHI Project, 2025).
@@
3.4 Model-based Spatial Analysis Methods
In addition to summary-based spatial descriptors, model-based approaches have been developed to analyze cellular spatial organization in histopathology and multiplex imaging data. These methods specify probabilistic or functional models under explicit structural assumptions, enabling formal statistical inference (Masotti et al., 2023; Seal et al., 2024; Mohammed et al., 2024).
Such approaches typically require model specification (e.g., parametric interaction functions or prior distributions), likelihood- or Bayesian-based estimation procedures, and tailored preprocessing steps. To contextualize the broader methodological landscape, several representative formulations are summarized below. These methods are not implemented within SASHIMI, which focuses on standardized spatial descriptors designed for scalable and interactive exploratory analysis.
3.4.1 Bayesian Hidden Potts Mixture Model
The Bayesian hidden Potts mixture model captures cellular heterogeneity and spatial dependence through a finite mixture model with Potts-type spatial regularization. Each spatial location $(x_{i},y_{i})$ is associated with a latent class label $m_{i}\in\{1,\dots,K\}$. Conditional on class membership, cell-level feature measurements are modeled via a finite mixture likelihood
$p(\boldsymbol{z}\mid\boldsymbol{m},\boldsymbol{\theta})=\prod_{i=1}^{n}f(z_{i}\mid\theta_{m_{i}}),$
where $z_{i}$ denotes observed features at location $i$, $\theta_{m_{i}}$ are class-specific parameters, and $\boldsymbol{\theta}=\{\theta_{k}\}_{k=1}^{K}$. Spatial dependence among neighboring latent labels is modeled using a Potts prior
$p(\boldsymbol{m}\mid\beta)\propto\exp\!\left(\beta\sum_{i\sim j}\mathbb{I}(m_{i}=m_{j})\right),$
where $i\sim j$ denotes neighboring locations and $\beta$ controls the strength of spatial smoothing. Bayesian inference is conducted under the joint posterior
$p(\boldsymbol{m},\boldsymbol{\theta},\beta\mid\boldsymbol{z})\propto p(\boldsymbol{z}\mid\boldsymbol{m},\boldsymbol{\theta})\,p(\boldsymbol{m}\mid\beta)\,p(\boldsymbol{\theta})\,p(\beta),$
allowing estimation of spatially structured latent classes and interaction strength (Li et al., 2019c).
3.4.2 Bayesian Mark Interaction Model
A Bayesian mark interaction model specifies a Gibbs distribution for the mark configuration conditional on observed spatial locations. Given the marked point pattern $\{(x_{i},y_{i},m_{i})\}_{i=1}^{n}$, the marks are treated as random variables conditional on fixed locations. The model is defined through a local energy function
$V(\boldsymbol{m}\mid\boldsymbol{\omega},\boldsymbol{\Theta},\lambda)=\sum_{q}\omega_{q}\sum_{i}\mathbb{I}(m_{i}=q)+\sum_{q,q^{\prime}}\theta_{qq^{\prime}}\sum_{i\sim j}e^{-\lambda d_{ij}}\mathbb{I}(m_{i}=q,m_{j}=q^{\prime}),$
where $\boldsymbol{\omega}=\{\omega_{q}\}$ are first-order intensity parameters, $\boldsymbol{\Theta}=\{\theta_{qq^{\prime}}\}$ are second-order interaction parameters indexed by cell-type pairs, and $\lambda$ controls exponential decay of interaction strength with distance $d_{ij}$. The corresponding likelihood takes the Gibbs form
$p(\boldsymbol{m}\mid\boldsymbol{\omega},\boldsymbol{\Theta},\lambda)=\frac{\exp\!\left[-V(\boldsymbol{m}\mid\boldsymbol{\omega},\boldsymbol{\Theta},\lambda)\right]}{Z(\boldsymbol{\omega},\boldsymbol{\Theta},\lambda)},$
where $Z(\cdot)$ is an intractable normalizing constant. Under a Bayesian framework, prior distributions are specified for $\boldsymbol{\omega}$, $\boldsymbol{\Theta}$, and $\lambda$, and inference is performed on their posterior distribution (Li et al., 2019a).
3.4.3 DIMPLE: Distance-matrix–based Modeling Approaches
DIMPLE summarizes inter-type spatial relationships through empirical distance distributions. For each pair of cell types $(p,q)$, inter-point distances $d_{ij}$ with $(x_{i},y_{i})\in X_{p}$ and $(x_{j},y_{j})\in X_{q}$ are computed and aggregated into a distribution $F_{pq}(r)$.
For two samples $a$ and $b$, differences between their interaction patterns are quantified using information-theoretic divergences. The Kullback–Leibler divergence is defined as
$\mathrm{KLD}\!\left(F_{pq}^{(a)}\,\|\,F_{pq}^{(b)}\right)=\sum_{r}F_{pq}^{(a)}(r)\log\frac{F_{pq}^{(a)}(r)}{F_{pq}^{(b)}(r)},$
and the symmetric Jensen–Shannon distance is
$\mathrm{JSD}\!\left(F_{pq}^{(a)},F_{pq}^{(b)}\right)=\sqrt{\frac{1}{2}\mathrm{KLD}\!\left(F_{pq}^{(a)}\,\|\,M\right)+\frac{1}{2}\mathrm{KLD}\!\left(F_{pq}^{(b)}\,\|\,M\right)},\quad M=\frac{1}{2}\!\left(F_{pq}^{(a)}+F_{pq}^{(b)}\right).$
For multiple samples, these divergences are assembled into sample-by-sample distance matrices for each interaction type $(p,q)$, and statistical inference is performed on these matrices to assess associations with group labels or clinical outcomes (Masotti et al., 2023).
3.4.4 SpaceANOVA: Point Process and Functional Modeling Approaches
SpaceANOVA models inter-type spatial dependence through distance-dependent co-occurrence functions combined with functional ANOVA. For two cell types $p$ and $q$, a co-occurrence function $C_{pq}(r)$ summarizes interaction intensity between $X_{p}$ and $X_{q}$ at spatial scale $r$, estimated from inter-point distances $d_{ij}$. For multiple samples indexed by $k$, the co-occurrence function is treated as a functional response:
$C_{pq}^{(k)}(r)=\mu_{pq}(r)+\alpha_{pq,g(k)}(r)+\varepsilon_{k}(r),$
where $\mu_{pq}(r)$ is a baseline interaction function, $\alpha_{pq,g(k)}(r)$ denotes group-specific functional effects, and $\varepsilon_{k}(r)$ is a stochastic error process. Functional ANOVA is used to test for differences in spatial interaction patterns across groups while accounting for spatial scale (Seal et al., 2024).